# EllipSect: a surface brightness analysis tool for GALFIT 3

### v1

@@ 0.6.1 Installing EllipSect
1pip install ellipse
Step-by-step installation
The following libraries are required to install EllipSect (v3.4.4) along with Python >=>=> = v3.8:
  • numpy >=>=> = v1.20.3
  • astropy >=>=> = v5.1
  • scipy >=>=> = v1.5.2
  • matplotlib >=>=> = v3.5.2
  • mgefit >=>=> = v5.0.13
  • sphinx >=>=> = v3.2.1
@@
1cd EllipSect
2python setup.py install
@@
1python -m pytest
Although GALFIT is not strictly required, it will require to create the model components and sigma image. Ensure you can call GALFIT from the command line. Otherwise, the automated tests will fail.
0.6.2 Running EllipSect
@@
1EllipSect GALFIT.01
Note: EllipSect takes the GALFIT output files (GALFIT.XX) where XX is the output file number you want to plot.
@@
1 EllipSect GALFIT.01 --phot
EllipSect has the option to compute the sky to compare with or provide to GALFIT. As explained above, the program uses two methods to compute it: The gradient method and the random box method. To compute the gradient method:
1 EllipSect GALFIT.01 -gsky
@@
1 EllipSect GALFIT.01 -rsky
@@
1EllipSect --help
Running EllipSect as a script
@@
1 from ellipsect.inout.read import ArgParsing
2 from ellipsect.sectors.sect import SectorsGalfit
3
4 args=[’GALFIT.01’,’--logx’, ’--phot’,’--noplot’]
5 parser_args = ArgParsing(args)
6 photapi = SectorsGalfit(parser_args)
7
8 print("Akaike Criterion: ", photapi.AICrit)
9 print("Bulge to Total: ", photapi.BulgeToTotal)

### v2

@@ 7.1.1 Quick installation
pip install ellipse
7.1.2 Step-by-step installation
The following libraries are required to install EllipSect (v3.4.4) along with Python >=>= v3.8 (python2009):
  • numpy >=>= v1.20.3 (Numpy2020)
  • astropy >=>= v5.1 (astropy:2013; astropy:2018)
  • scipy >=>= v1.5.2 (Scipy2020)
  • matplotlib >=>= v3.5.2 (matplotlib2007)
  • mgefit >=>= v5.0.13 (cappellari02)
  • sphinx >=>= v3.2.1 (sphinx2021)
@@ 7.1.2 Step-by-step installation
cd EllipSect
python setup.py install
@@
python -m pytest
Although GALFIT is not strictly required, it will require to create the model components and sigma image. Ensure you can call GALFIT from the command prompt. Otherwise, the automated tests will fail.
7.2 Running EllipSect
@@
EllipSect GALFIT.01
Note: EllipSect takes the GALFIT output files (GALFIT.nn) where nn is the output file number you want to plot.
@@
EllipSect GALFIT.01 phot
EllipSect has the option to compute the sky level to compare with or to provide to GALFIT. As explained above, the program uses two methods to compute it: The gradient method and the random box method. To compute the gradient method:
EllipSect GALFIT.01 -gsky
@@
EllipSect GALFIT.01 -rsky
@@
EllipSect help
7.2.1 Running EllipSect as a script
@@
from ellipsect.inout.read import ArgParsing
from ellipsect.sectors.sect import SectorsGalfit
args=[’GALFIT.01’,’–logx’, ’–phot’,’–noplot’]
parser_args = ArgParsing(args)
photapi = SectorsGalfit(parser_args)
print(”Akaike Criterion: ”, photapi.AICrit)
print(”Bulge to Total: ”, photapi.BulgeToTotal)